# From Taylor Series to Fourier Synthesis: The Periodic Linear Unit

### v1

@@ 2.1 Choices Made
The scaling factors are important hyperparemeters. $\beta$ provides balancing between the sine wave and the linear component. $\beta=1.0$ is a good default value, as it strikes a delicate balance between the periodic gradient ”deadzone” that begins occuring every $\pi$ cycles if it is set much higher, and the completely non-periodic simple linear unit obtained from setting it to 0.00.00.0. $\alpha$ provides the means for modifying the period of the activation.
@@
    While the model is given freedom to increase the oscillation of the activation through changing the $\beta$, there is a point at which such oscillations become a major issue. A rapidly oscillating activation deteorates the loss landscape significantly with its severe non-monotonicity, and can lead to numerical stability issues or training runs that have to be stopped midway. Having the scaling term be formulated as $\frac{\beta}{1+\lvert\beta\rvert}$ however, means that the activation discourages runaway activations like this. It discourages the model from introducing wildly fluctuating activations, as you have to have $\beta$ become very large to obtain such results. The maximum oscillation which can be introduced by the model is $x+\sin(x)$ at $\beta\to\infty$, effectively capping the oscillating behavior at a well-behaved range.
@@
[…]
@@
    An activation in the same close family of PLU is Snake, which is formulated as $x+\frac{1}{a}\sin^{2}(ax)$. As we will notice immediately, if the learnable parameter $a$ needs to cross zero during training (e.g., to flip the ”phase” of the non-linearity), the model encounters a 1.0/0.01.0/0.01.0 / 0.0 singularity. This results in NaNs and a crashed training run, and is usually fixed by a custom autograd function. PLU however does not require this: Since $\lvert\beta\rvert$ is always non-negative (0$\geq$ 0≥ 0), the denominator $1+\lvert\beta\rvert$ is always greater than or equal to 1. Hence, there is no possibility of division by zero. The function is perfectly defined and stable for all possible real values of $\beta$.
Another distinct feature is the $\lvert\alpha\rvert$ term. Why is there an absolute value function applied to $\alpha$? This was chosen based on the consideration of phase-flipping the activation function. One of the features of PLU is the ability for the model to flip the phase of the oscillation if necessary, but a crucial consideration when giving the model such an ability is then the various avenues with which it can achieve this. Importantly, notice that without the absolute value on the $\alpha$, the model has two ways of flipping the phase: either by letting $\beta$, the periodicity strength parameter cross 0.00.00.0, or by having $\alpha$ cross zero. Imagine that a specific gradient step pushes the model in the direction of flipping the phase. What would happen to the gradients at both those parameters? The gradient step might push both $\alpha$ and $\beta$ across 0.00.00.0. But if this step is taken, then the phase flip is essentially cancelled. We flip it twice and end up with the same phase. This is not good for the loss landscape, and so we must choose one of these mechanisms and stick to using it for flipping the oscillatory behavior. It is then much more natural to use $\beta$ for this purpose, as crossing 0.00.00.0 with $\beta$ does not introduce any wild changes with the activation, while $\alpha$ crossing 0.00.00.0 pushes the scale of the entire activation horizontally and massively, letting the period of the oscillation grow to infinity, before coming back to a reasonable range once again with flipped phase.
@@ 2.2 The Untapped Potential of Non-Monotonicity
At certain values of $\alpha$ and $\beta$, namely when $1+\frac{\beta}{1+\lvert\beta\rvert}\lvert\alpha\rvert\cos(\lvert\alpha\rvert x)<0$ for some $x$, PLU turns into a non-monotonically increasing function. This is not the case for the default $(\alpha,\beta)$ values of (1.0,1.0)(1.0,1.0)( 1.0 , 1.0 ), but in a parametric setting, it is possible for the model to learn this behavior.
To control the amount of non-monotonicity allowed, as we have discussed already in the point above about capping the oscillating behavior to a well-behaved range, this is done by making non-monotonicity of an excessive degree harder to obtain by the model, as a large $\beta$ of around 10.010.010.0 is necessary to even approach a 0.90.90.9 influence from the sine wave relative to the residual $x$. Optimizers tend to prefer moderately sized parameters, and SGD steps are small, meaning this is a subtle push that encourages the model to utilize moderate amounts of the periodic component and focus on improving the model elsewhere.
@@
By taking a closer look at the potential differences, we can quickly ascertain that compared to monotonic activations like ReLU when $x>0$ or Snake, the training dynamics will be entirely different, as in those cases, the monotonic activations basically act as transparent layers that don’t affect the SGD steps per-se. At each step, monotonic activations don’t fundamentally alter the direction of the gradient since $dA/dx$ (where $A$ is the activation) is always positive and by the chain rule, applying said partial derivative will never flip the sign; thus, these activations end up acting more like a dynamic learning rate than a complex function shaper.
@@ 2.3 Asymptotic Behavior
  • As $\beta\to+\infty$, the term approaches +1+1+ 1, and the activation becomes $x+\sin(\alpha x)$.
  • As $\beta\to-\infty$, the term approaches 1-1- 1, and the activation becomes $x-\sin(\alpha x)$.
  • If $\beta=1$, the term is 1/2=0.51/2=0.51 / 2 = 0.5. The activation becomes $x+0.5\sin(\alpha x)$, where the derivative at $x=0$ is 1.51.51.5.
  • If $\beta=-1$, the term is 1/2=0.5-1/2=-0.5- 1 / 2 = - 0.5. The activation becomes $x-0.5\sin(\alpha x)$, where the derivative at $x=0$ is 0.50.50.5.
@@ 4 Repulsive Reparameterization
The key in solving the problem of helping the optimizer learn to utilize the complex activation function is the often reinforced observation about the path of least resistance. If there are two directions which the model weights can go in, one that immediately raises the loss immensely but with the unknowable potential of a lower loss beyond the cliff, and another that promises to lower the loss slightly, the gradient step will overwhelmingly be in the direction of lowering the loss. Mechanisms such as the RMSProp [2] or Adam [3] optimizers attempt to fix this, but even they will not let a model take a step that appears to increse the loss ten-fold.
Thus, the key in forcing the optimizer to use PLU to its full effect turns out to be in purposefully making the increse in loss from turning it into a linear function immense. While previously, the optimizer might have, at the very first step, seen the two paths and decided that the loss can be lowered by reducing the PLU to a linear function, we now seek for the optimizer to instead see a massive towering cliff in front of the path towards a more linear PLU, while the path away is unimpeded.
Specifically, the problematic values for both $\alpha$ and $\beta$ are when they are 0.00.00.0. So, in effect we hope that the penalty in the gradients for lowering loss grows more and more as the optimizer pushes PLU towards having near-zero $\alpha$ and $\beta$’s, thus discouraging the optimizer from collapsing the activation into a linear identity.
@@ 5 The Periodic Linear Unit
As the optimizer attempts to push $\alpha$ towards 0.00.00.0, the $\rho_{\alpha}/\alpha$ term grows quickly, exploding to infinity and thus making the period of the sine component become small at an alarming rate. While many real world data points have the tendency to have repeating patterns in them, there is none that will benefit in any way from a sine component with a period that is quickly growing infinitely small. Thus, the optimizer is actively pushed back against its desires to push $\alpha$ towards 0.00.00.0 by an explosion of loss when taking even a small step in this direction.
As the optimizer attempts to push $\beta$ towards 0.00.00.0, the $\rho_{\beta}/\beta$ term causes $\beta_{\text{eff}}$ to explode towards positive or negative infinity. Consequently, the amplitude scaling term $\beta_{\text{eff}}/(1+\lvert\beta_{\text{eff}}\rvert)$ rapidly approaches its maximum absolute value of 1. The periodic component becomes ”stuck” at full amplitude (±$\pm\sin(\lvert\alpha_{\text{eff}}\rvert*x)$), and the optimizer thus loses all ability to reduce its influence. To regain control and continue minimizing the loss, it is then forced to push $\beta$ away from the zero-crossing, thus avoiding the linear identity function.
The most crucial part of this reparameterization is that for a sufficiently large $\alpha$ and $\beta$, the reparameterization does not occur. The original $\alpha$ and $\beta$ are simply passed through, as both $\rho_{p}/p$ terms approach 0.00.00.0 as $p=(\alpha\text{ or }\beta)$ grows. This is the mathematical formulation of our ”gradient wall” idea. In the end, all that is affected is the asympototic behavior at $\alpha=0.0$ and $\beta=0.0$, and the full expressive behavior of PLU is left fully intact. We term this mechanism Repulsive Reparameterization.
@@ 5.1 Choice of $\rho_{\alpha}$ and $\rho_{\beta}$
Thus, the reparametrization scheme in effect limits the effective range of $P$ to $P_{\text{eff}}\geq P_{0}$, or more intuitively $P_{\text{eff}}\geq\sqrt{\rho_{P}}$.
@@
$\alpha$ is the period multiplier, and a larger value produces a PLU which oscillates faster. A lower value on the other hand creates longer oscillations. In this case then we should set $\rho_{\alpha}$ based on the lower bound of how wide of an activation oscillation we want to allow. A $\rho_{\alpha}$ of 5.05.05.0, as an example, will allow any $\alpha_{\text{eff}}\geq\sqrt{5.0}\approx 2.236$. An overly large $\rho_{\alpha}$ will force the model into utilizing a PLU activation with a fast oscillation by enforcing a minimum frequency on the sin component in this way, the effects of which are explored in our ’High-Frequency Prior’ experiment (Section 6.2, “The ’Chaotic Initialization’ Paradigm”). The effects of this are interesting to say the least, and will be explored later, as they indicate a profound shift in how neural networks can learn.
$\beta$ is then the soft-bounded strength of the sine component. In the previous section it was mentioned that a $\beta$ value of 0.00.00.0 corresponds to a linear function, and as it is increased, slowly, more of the sine component is introduced. $\rho_{\beta}$ then provides a lower bound for $\beta$, forcing the model to never be able to go below a specific sine wave contribution fraction. For example, a $\rho_{\beta}$ of 0.150.150.15 will allow only $\beta_{\text{eff}}\geq\sqrt{0.15}\approx 0.387$, or in direct terms a minimum multiplier on the sin component of $\beta_{\text{eff}}/(1+\lvert\beta_{\text{eff}}\rvert)\approx 0.279$.
@@ 6.1 Superior Convergence and Expressivity in a Standard MLP
[…]
@@ 6.2 The “Chaotic Initialization” Paradigm
[…]

### v2

@@ 2.1 Choices Made
The scaling factors are important hyperparameters. $\beta$ provides balancing between the sine wave and the linear component. $\beta=1.0$ is a good default value, as it strikes a delicate balance between the periodic gradient ”deadzone” that begins occurring every $\pi$ cycles if it is set much higher, and the completely non-periodic simple linear unit obtained from setting it to $0.0$. $\alpha$ provides the means for modifying the period of the activation.
@@
    While the model is given freedom to increase the oscillation of the activation through changing the $\beta$, there is a point at which such oscillations become a major issue. A rapidly oscillating activation deteriorates the loss landscape significantly with its severe non-monotonicity, and can lead to numerical stability issues or training runs that have to be stopped midway. Having the scaling term be formulated as $\frac{\beta}{1+\lvert\beta\rvert}$ however, means that the activation discourages runaway activations like this. It discourages the model from introducing wildly fluctuating activations, as you have to have $\beta$ become very large to obtain such results. The maximum oscillation which can be introduced by the model is $x+\sin(x)$ at $\beta\to\infty$, effectively capping the oscillating behavior at a well-behaved range.
@@
    The activation can provide a very good selection of shapes with $\alpha$ and $\beta$ both being just around the $3.0<\text{value}<3.0$ range, which is very good considering optimizers tend to prefer weights in that general range as well. You can get a flipped phase activation by flipping the sign on $\beta$, you can get non-monotonicity with the right combinations, and more.
@@
    An activation in the same close family of PLU is Snake, which is formulated as $x+\frac{1}{a}\sin^{2}(ax)$. As we will notice immediately, if the learnable parameter $a$ needs to cross zero during training (e.g., to flip the ”phase” of the non-linearity), the model encounters a $1.0/0.0$ singularity. This results in NaNs and a crashed training run, and is usually fixed by a custom autograd function. PLU however does not require this: Since $\lvert\beta\rvert$ is always non-negative ($\geq 0$), the denominator $1+\lvert\beta\rvert$ is always greater than or equal to 1. Hence, there is no possibility of division by zero. The function is perfectly defined and stable for all possible real values of $\beta$.
Another distinct feature is the $\lvert\alpha\rvert$ term. Why is there an absolute value function applied to $\alpha$? This was chosen based on the consideration of phase-flipping the activation function. One of the features of PLU is the ability for the model to flip the phase of the oscillation if necessary, but a crucial consideration when giving the model such an ability is then the various avenues with which it can achieve this. Importantly, notice that without the absolute value on the $\alpha$, the model has two ways of flipping the phase: either by letting $\beta$, the periodicity strength parameter cross $0.0$, or by having $\alpha$ cross zero. Imagine that a specific gradient step pushes the model in the direction of flipping the phase. What would happen to the gradients at both those parameters? The gradient step might push both $\alpha$ and $\beta$ across $0.0$. But if this step is taken, then the phase flip is essentially cancelled. We flip it twice and end up with the same phase. This is not good for the loss landscape, and so we must choose one of these mechanisms and stick to using it for flipping the oscillatory behavior. It is then much more natural to use $\beta$ for this purpose, as crossing $0.0$ with $\beta$ does not introduce any wild changes with the activation, while $\alpha$ crossing $0.0$ pushes the scale of the entire activation horizontally and massively, letting the period of the oscillation grow to infinity, before coming back to a reasonable range once again with flipped phase.
@@ 2.2 The Untapped Potential of Non-Monotonicity
At certain values of $\alpha$ and $\beta$, namely when $1+\frac{\beta}{1+\lvert\beta\rvert}\lvert\alpha\rvert\cos(\lvert\alpha\rvert x)<0$ for some $x$, PLU turns into a non-monotonically increasing function. This is not the case for the default $(\alpha,\beta)$ values of $(1.0,1.0)$, but in a parametric setting, it is possible for the model to learn this behavior.
To control the amount of non-monotonicity allowed, as we have discussed already in the point above about capping the oscillating behavior to a well-behaved range, this is done by making non-monotonicity of an excessive degree harder to obtain by the model, as a large $\beta$ of around $10.0$ is necessary to even approach a $0.9$ influence from the sine wave relative to the residual $x$. Optimizers tend to prefer moderately sized parameters, and SGD steps are small, meaning this is a subtle push that encourages the model to utilize moderate amounts of the periodic component and focus on improving the model elsewhere.
@@
By taking a closer look at the potential differences, we can quickly ascertain that compared to monotonic activations like ReLU when $x>0$ or Snake, the training dynamics will be entirely different, as in those cases, the monotonic activations basically act as transparent layers that don’t affect the SGD steps per-se. At each step, monotonic activations don’t fundamentally alter the direction of the gradient since $dA/dx$ (where AA is the activation) is always positive and by the chain rule, applying said partial derivative will never flip the sign; thus, these activations end up acting more like a dynamic learning rate than a complex function shaper.
@@ 2.3 Asymptotic Behavior
  • As $\beta\to+\infty$, the term approaches +1+1, and the activation becomes $x+\sin(\alpha x)$.
  • As $\beta\to$-$\infty$, the term approaches 1-1, and the activation becomes $x-\sin(\alpha x)$.
  • If $\beta=1$, the term is $1/2=0.5$. The activation becomes $x+0.5\sin(\alpha x)$, where the derivative at $x=0$ is $1.5$.
  • If $\beta=-1$, the term is 1/2=0.5-1/2=-0.5. The activation becomes $x-0.5\sin(\alpha x)$, where the derivative at $x=0$ is $0.5$.
@@ 4 Repulsive Reparameterization
The key in solving the problem of helping the optimizer learn to utilize the complex activation function is the often reinforced observation about the path of least resistance. If there are two directions which the model weights can go in, one that immediately raises the loss immensely but with the unknowable potential of a lower loss beyond the cliff, and another that promises to lower the loss slightly, the gradient step will overwhelmingly be in the direction of lowering the loss. Mechanisms such as the RMSProp [2] or Adam [3] optimizers attempt to fix this, but even they will not let a model take a step that appears to increase the loss ten-fold.
Thus, the key in forcing the optimizer to use PLU to its full effect turns out to be in purposefully making the increase in loss from turning it into a linear function immense. While previously, the optimizer might have, at the very first step, seen the two paths and decided that the loss can be lowered by reducing the PLU to a linear function, we now seek for the optimizer to instead see a massive towering cliff in front of the path towards a more linear PLU, while the path away is unimpeded.
Specifically, the problematic values for both $\alpha$ and $\beta$ are when they are $0.0$. So, in effect we hope that the penalty in the gradients for lowering loss grows more and more as the optimizer pushes PLU towards having near-zero $\alpha$ and $\beta$’s, thus discouraging the optimizer from collapsing the activation into a linear identity.
@@ 5 The Periodic Linear Unit
As the optimizer attempts to push $\alpha$ towards $0.0$, the $\rho_{\alpha}/\alpha$ term grows quickly, exploding to infinity and thus making the period of the sine component become small at an alarming rate. While many real world data points have the tendency to have repeating patterns in them, there is none that will benefit in any way from a sine component with a period that is quickly growing infinitely small. Thus, the optimizer is actively pushed back against its desires to push $\alpha$ towards $0.0$ by an explosion of loss when taking even a small step in this direction.
[…]
The most crucial part of this reparameterization is that for a sufficiently large $\alpha$ and $\beta$, the reparameterization does not occur. The original $\alpha$ and $\beta$ are simply passed through, as both $\rho_{p}/p$ terms approach $0.0$ as $p=(\alpha\text{ or }\beta)$ grows. This is the mathematical formulation of our ”gradient wall” idea. In the end, all that is affected is the asymptotic behavior at $\alpha=0.0$ and $\beta=0.0$, and the full expressive behavior of PLU is left fully intact. We term this mechanism Repulsive Reparameterization.
@@ 5.1 Choice of $\rho_{\alpha}$ and $\rho_{\beta}$
Thus, the reparametrization scheme in effect limits the effective range of $P$ to $P_{\text{eff}}\geq 2P_{0}$, or more intuitively $P_{\text{eff}}\geq 2\sqrt{\rho_{P}}$.
@@
$\alpha$ is the period multiplier, and a larger value produces a PLU which oscillates faster. A lower value on the other hand creates longer oscillations. In this case then we should set $\rho_{\alpha}$ based on the lower bound of how wide of an activation oscillation we want to allow. A $\rho_{\alpha}$ of $5.0$, as an example, will allow any $\alpha_{\text{eff}}\geq\sqrt{5.0}\approx 2.236$. An overly large $\rho_{\alpha}$ will force the model into utilizing a PLU activation with a fast oscillation by enforcing a minimum frequency on the sin component in this way, the effects of which are explored in our ’High-Frequency Prior’ experiment (Section 6.2, “The ’Chaotic Initialization’ Paradigm”). The effects of this are interesting to say the least, and will be explored later, as they indicate a profound shift in how neural networks can learn.
$\beta$ is then the soft-bounded strength of the sine component. In the previous section it was mentioned that a $\beta$ value of $0.0$ corresponds to a linear function, and as it is increased, slowly, more of the sine component is introduced. $\rho_{\beta}$ then provides a lower bound for $\beta$, forcing the model to never be able to go below a specific sine wave contribution fraction. For example, a $\rho_{\beta}$ of $0.15$ will allow only $\beta_{\text{eff}}\geq\sqrt{0.15}\approx 0.387$, or in direct terms a minimum multiplier on the sin component of $\beta_{\text{eff}}/(1+\lvert\beta_{\text{eff}}\rvert)\approx 0.279$.
@@ 6.1 Superior Convergence and Expressivity in a Standard MLP
Our first experiment utilizes a standard MLP architecture with two hidden layers of 8 neurons each (2-8-8-1). We initialize PLU with moderate repulsion ($\text{init\_rho\_alpha}=5.0$, $\text{init\_rho\_beta}=0.15$) (The Adam optimizer begins with these initial values of $\rho_{\alpha}$ and $\rho_{\beta}$ and is allowed to modify them).
@@ 6.2 The “Chaotic Initialization” Paradigm
To explore the effect of the repulsive reparameterization, we conduct a second experiment with the same 8-neuron 2-8-8-1 MLP but initialize PLU with extreme repulsion terms ($\text{init\_rho\_alpha}=15.0$, $\text{init\_rho\_beta}=0.5$). This forces the effective $\alpha$ and $\beta$ to be large from the outset, inducing high-frequency oscillations. Specifically, the earlier configurations for $\rho$ sets a lower-bound for the optimizer on $\alpha$ and $\beta$ of $\alpha_{\text{eff}}\geq 2.236$ and $\beta_{\text{eff}}\geq 0.387$, while with the new configurations, this lower floor is lifted all the way to $\alpha_{\text{eff}}\geq 3.873$ and $\beta_{\text{eff}}\geq 0.707$. The effect is that the contribution from the sine wave is forced to be very high throughout the run, and the oscillation period of that sine wave is also kept incredibly short.